# Probing the band structure of the strongly correlated antiferromagnet NiPS3 across its phase transition

### v1

@@ II Results and Discussion
This reasonable interpretation leaves us with the question of the origin of the flat feature in ARPES at eV, that is not captured by DFT+U. It appears as a shoulder in the raw ARPES data at energies above the valence band maximum and is nearly identical above and below (Suppl. Fig S11). It, moreover, exhibits the identical photon energy () dependence as the valence band maximum (Suppl. Fig. S5, S11). To substantiate its discussion, we compare with the data of MnPS3 Strasdas et al. (2023) (Fig. 5), where all observed features in the ARPES data could be identified with bands of the DFT+U calculation. Figure 5a-f displays the direct comparison using DFT+U data within the respective antiferromagnetic ground state. The figure includes the partial density of states (PDOS) (Fig. 5a,f) where the DOS is projected to the S and P atoms (left) or the eg (3d) and t2g (3d) orbitals of Ni and Mn (right), respectively, using the octahedral coordinates with along a Ni-S bond (Fig. 5b,e). Moreover, an additional distinction between minority and majority spins within an individual atom has been performed for the Ni 3d levels. Finally, the ARPES and DFT+U data of NiPS3 are shifted upwards by 0.4 eV with respect to their to align the low energy bands of the two materials. Hence, is the Fermi level of MnPS3 only.
@@
We find large similarities between NiPS3 and MnPS3. For example, the upper valence band of S 3p/Ni 3d character, respectively S 3p/Mn 3d character is nearly identical in strength of dispersion and orbital contributions. Moreover, the lowest energy bands ( eV) of mixed S 3p/P 3p character are very similar including the relative contribution of S and P orbitals. The energy range in between is, however, different. Firstly, there is a full spin polarization of the t2g orbitals for MnPS3 that is absent for NiPS3. Related, the mixed 3d/S 3p bands are more confined in energy for MnPS3. In particular, the two bands dominated by Mn 3d orbitals ( eV ) are the ones that are most strongly changed across Strasdas et al. (2023). In contrast, for NiPS3, all bands with eV have strong, often dominating contributions from Ni 3d orbitals (Suppl. section S4). This difference is in line with the 3d level occupation (see below). Most importantly, the experimental data is very reasonably fitted by DFT+U for both data sets except for the barely dispersing feature at eV for NiPS3 (highlighted by the yellow box in Fig. 5c). The correspondence between ARPES and DFT+U is also reasonable for FePS3, albeit a stretch of DFT+U data by 7 % was required for a quantitative match Pestka et al. (2024).
@@
Using the octahedral coordinates along the bond directions to assign the 3d levels leads to a nice visualization of the expected 3d level occupation for the two materials (Fig. 5g-l). For Mn2+ 3d5, a complete spin polarization of the 3d levels is obtained (Fig. 5k) as expected from Hund’s rule (Fig. 5l). In addition, the occupied eg levels are split into two parts due to the deviations from a perfect octahedral symmetry Figgis and Hitchman (2000); Trimarchi et al. (2018). On the other hand, Ni2+ 3d8 shows completely occupied t2g levels, while only the eg levels are close to fully spin-polarized (Fig. 5h). Hence, the occupation of bands is in line with the expected 3d level occupations and energy splittings. Moreover, for both materials, the bands that are changing most strongly across exhibit significant 3d contributions. This makes us confident that the band assignment by DFT+U is correct and the feature at eV is beyond the related approximations. It is very likely that 3d level multiplets are involved, maybe including ligand contributions Kang et al. (2020), as will be discussed in a separate publication.
@@ III Conclusions
We probed the band structure of NiPS3 above and below the Néel temperature using -ARPES on flakes exfoliated onto a Au surface. The ARPES data are compared with DFT+U calculations to identify the orbital character of bands. We identify an upwards band shift below by roughly 150 meV for a band close to . This band consists of mixed Ni 3d and S 3p orbitals that likely feature an antiferromagnetic Ni 3t2g-S 3p-Ni 3t2g superexchange path. Beyond the good agreement between DFT+U and ARPES that we also found for FePS3 Pestka et al. (2024) and MnPS3 Strasdas et al. (2023), we identified a shoulder in the ARPES data above the valence band maximum that could not be attributed to a band of the DFT+U data. This is in striking contrast to FePS3 and MnPS3. We assume that the additional structure is reminiscent of the known strong correlations in NiPS3.
@@
Sample preparation: We employ an optimized exfoliation method (scotch tape: Nitto ELP BT-150E-KL) to prepare samples with few-layered flakes of NiPS3 in a cleanroom environment. Commercially available Si/SiO2 (oxide thickness 90 nm) is used as substrate. It has been metal-coated at 370 K with Au (5 nm)/Ti (1 nm) in an ultrahigh vacuum e-beam evaporator. This improves the flake adhesion and renders the substrate conducting, which is necessary to prevent charging during ARPES Strasdas et al. (2023); Koitzsch et al. (2023); Velický et al. (2018). A mild O2 plasma ashing (power = 50 W, flux = 100 sccm) is done at room temperature for 25 s just before exfoliating NiPS3 flakes to efficiently remove organic molecules from the gold surface by oxidation. NiPS3 is exfoliated at 60∘C within few minutes after the plasma cleaning. This leads to flakes with diverse thickness, where the desired thickness and flake size are identified using commercial optical and atomic force microscopes Strasdas et al. (2023).
@@
ARPES and XPS experiments: -ARPES and XPS were performed at the NanoESCA beamline of the Elettra synchrotron radiation facility in Italy. An ultrahigh vacuum FOCUS NanoESCA photoemission electron microscope (PEEM) was used for the -space mapping mode operation working at a background pressure of mbar Schneider et al. (2012). To ensure clean surfaces, the samples are firstly annealed at 200∘C for a couple of hours prior to ARPES. The synchrotron allows to focus a wide range of photon energies on the sample with a beam spot of 5-10 m in diameter and total energy resolution (beamline and analyzer) of 50 meV. For our experiments, the photon beam impinged with an incident angle of with respect to the surface normal and with a azimuth relative to the direction of the crystal. The polarization was within the plane of incidence (p-polarized). The Fermi level has been determined by ARPES on the Au film of the substrate. The sample temperature is adjusted by changing the liquid He flow rate and measured by a Lake Shore DT-670E-BR Si diode placed at the sample holder.
ARPES curvature: We partially use ARPES curvature plots for better visibility of the bands Zhang et al. (2011). The curvature is calculated with respect to energy after smoothing the raw data, as described in the caption of Fig. 2c, via:
and are the first and second partial derivative with respect to energy of the smoothed photoelectron intensity , is the maximum value of the first partial derivative, while is a free parameter used to adjust the relative intensity of different bands for better visibility. Here, we use .
@@
Computational Details: The calculations were performed in the framework of density functional theory (DFT) using the generalized gradient approximation within the PBE flavor Perdew et al. (1996), as implemented in VASP software Kresse and Müller (1996). The ion–electron interactions were described by the projector augmented wave (PAW) method Holzwarth et al. (2001). Plane-wave basis cutoff and centered Monkhorst-Pack Monkhorst and Pack (1976) k-point grid were set to 550 eV and , respectively. A Gaussian smearing of 0.05 eV was employed for the Brillouin zone (BZ) integration. The interlayer vdW forces were treated within the Grimme scheme using D3 correction Grimme et al. (2010). All of the results were obtained using PBE+U method based on Dudarev’s approach Dudarev et al. (1998), with the effective on-site Hubbard U parameter (, where is fixed to eV) for the 3d orbitals. The position of the atoms and unit cell were fully optimized within the PBE+U approach. To simulate the disordered paramagnetic state, we used a supercell with spin-orbit interaction and randomly set the directions of the magnetic spin moments (in 3D) so that the net moment was zero. Here, we did not optimize the geometric structure, in order to capture only changes in electronic dispersion. Moreover, we used the widely employed band unfolding method to obtain an effective electronic structure Popescu and Zunger (2012).
@@
All data that support the findings of this study are available in zenodo at DOI 10.5281/zenodo.16162322.
@@
Details on sample preparation, experimental and computational methods, adaption of and including dependence of ARPES data, individual orbital projections of bands, additional charge density plots, ARPES and DFT band structure along , DFT+U band structure of paramagnetic phase and analysis of ARPES structure at eV.
@@ Acknowledgement
We acknowledge financial support from the German Research Foundation (DFG) via the project Mo 858/19-1 and Germany’s Excellence Strategy Cluster of Excellence: Matter and Light for Quantum Computing(ML4Q) EXC 2004/1 (project Nr. 390534769) from the German Ministry of Education and Research (Project 05K2022 -ioARPES). B.B. acknowledges the support of Humboldt Research Fellowship funded by the Alexander von Humboldt Foundation. M.B. acknowledges financial support from the University of Warsaw under the "Excellence Initiative - Research University" project. PL-Grid Polish Infrastructure for Supporting Computational Science in the European Research Space, and of the Interdisciplinary Center of Modeling (ICM), University of Warsaw are gratefully acknowledged. K.W. acknowledges support by the National Science Centre in Poland under the project 2024/55/B/ST3/03144. A.K.B. and E.L. were supported by the European Commission via the Marie-Sklodowska Curie action Phonsi (H2020-MSCA-ITN-642656).

### v2

@@ II Results and Discussion
This reasonable interpretation leaves us with the question of the origin of the flat feature in ARPES at eV, that is not captured by DFT+U. It appears as a shoulder in the raw ARPES data at energies above the valence band maximum and is nearly identical above and below (Suppl. Fig S12). It, moreover, exhibits the identical photon energy () dependence as the valence band maximum (Suppl. Fig. S5, S12). To substantiate its discussion, we compare with the data of MnPS3 Strasdas2023 (Fig. 5), where all observed features in the ARPES data could be identified with bands of the DFT+U calculation. Figure 5a-f displays the direct comparison using DFT+U data within the respective antiferromagnetic ground state. The figure includes the partial density of states (PDOS) (Fig. 5a,f) where the DOS is projected to the S and P atoms (left) or the eg (3d) and t2g (3d) orbitals of Ni and Mn (right), respectively, using the octahedral coordinates with along a Ni-S bond. Moreover, the PDOS of minority and majority spins is plotted separately. Finally, the ARPES and DFT+U data of NiPS3 are shifted upwards by 0.4 eV with respect to their to align the low energy bands of the two materials. Hence, , the Fermi level of MnPS3, is chosen as reference for for the energy scale.
@@
We find large similarities between NiPS3 and MnPS3. For example, the upper valence band of S 3p/Ni 3d character, respectively S 3p/Mn 3d character is nearly identical in strength of dispersion and orbital contributions. Moreover, the lowest energy bands ( eV) of mixed S 3p/P 3p character are very similar including the relative contribution of S and P orbitals. The energy range in between is, however, different. Firstly, there is a full spin polarization of the t2g orbitals for MnPS3 that is absent for NiPS3. Related, the mixed 3d/S 3p bands are more confined in energy for MnPS3. In particular, the two bands dominated by Mn 3d orbitals ( eV ) are the ones that are most strongly changed across Strasdas2023. In contrast, for NiPS3, all bands with eV have strong, often dominating contributions from Ni 3d orbitals (Suppl. section S5). This difference is in line with the 3d level occupation (see below). Most importantly, the experimental data is very reasonably fitted by DFT+U for both data sets except for the barely dispersing feature at eV for NiPS3 (highlighted by the yellow box in Fig. 5c). The correspondence between ARPES and DFT+U is also reasonable for FePS3, albeit a stretch of DFT+U data by 7 % was required for a quantitative match Pestka2024.
To corroborate that the extra feature (yellow box, Fig. 5c) is not captured by DFT-type calculations, we also performed calculation with other approaches to treat the exchange-correlation energies. We employed HSE while varying the percentage of the Hartree-Fock exchange contribution as well as G0W0 and EVGW0 based on a starting PBE band structure (Suppl. section S13). The results feature various shifts of the Ni 3d PDOS, but cannot explain the appearance of an extra feature above the bunch of Ni related bands as observed in ARPES.
@@
Finally, we visualize the expected 3d level occupation for NiPS3 and MnPS3 by using the octahedral coordinates along the bond directions to assign the 3d level orbitals (Fig. 5g-l). For Mn2+ 3d5, a complete spin polarization of the 3d levels is obtained (Fig. 5k) as expected from Hund’s rule (Fig. 5l). In addition, the occupied eg levels are split into two parts due to the deviations from a perfect octahedral symmetry Figgis2000; Trimarchi2018. On the other hand, Ni2+ 3d8 shows completely occupied t2g levels, while only the eg levels are close to fully spin-polarized (Fig. 5h). Hence, the occupation of bands is in line with the expected 3d level occupations and energy splittings. Moreover, for both materials, the bands that are changing most strongly across exhibit significant 3d contributions. This makes us confident that the band assignment by DFT+U is correct and the feature at eV is beyond the related approximations. It is very likely that 3d level multiplets are involved, maybe including ligand contributions Kang2020, as will be discussed in a separate publication.
@@ III Conclusions
We probed the band structure of NiPS3 above and below the Néel temperature using -ARPES on flakes exfoliated onto a Au surface. The ARPES data are compared with DFT+U calculations to identify the orbital character of bands. We identify an upwards band shift below by roughly 150 meV for a band close to . This band consists of mixed Ni 3d and S 3p orbitals that likely feature an antiferromagnetic Ni 3t2g-S 3p-Ni 3t2g superexchange path. Beyond the good agreement between DFT+U and ARPES that we also found for FePS3 Pestka2024 and MnPS3 Strasdas2023, we identified a shoulder in the ARPES data above the valence band maximum that could not be attributed to a band of the DFT+U data. This is in striking contrast to FePS3 and MnPS3. We assume that the additional structure is reminiscent of the known strong correlations in NiPS3.
Note added in proof: A very recent publication also shows -dependent ARPES data of NiPS3 revealing similar band structure features in reasonable correspondance with a similarly selected eV Cao2025.
@@
Sample preparation: We employ an optimized exfoliation method (scotch tape: Nitto ELP BT-150E-KL) to prepare samples with few-layered flakes of NiPS3 in a cleanroom environment. Commercially available Si/SiO2 (oxide thickness 90 nm) is used as substrate. It has been metal-coated at 370 K with Au (5 nm)/Ti (1 nm) in an ultrahigh vacuum e-beam evaporator. This improves the flake adhesion and renders the substrate conducting, which is necessary to prevent charging during ARPES Strasdas2023; Koitzsch2023; Velicky2018. A mild O2 plasma ashing (power = 50 W, flux = 100 sccm) is done at room temperature for 25 s just before exfoliating NiPS3 flakes to efficiently remove organic molecules from the gold surface by oxidation. NiPS3 is exfoliated at 60∘C within few minutes after the plasma cleaning. This leads to flakes with diverse thickness, where the desired thickness and flake size are identified using commercial optical and atomic force microscopes Strasdas2023.
@@
ARPES and XPS experiments: -ARPES and XPS were performed at the NanoESCA beamline of the Elettra synchrotron radiation facility in Italy. An ultrahigh vacuum FOCUS NanoESCA photoemission electron microscope (PEEM) was used for the -space mapping mode operation working at a background pressure of mbar Schneider2012. To ensure clean surfaces, the samples are firstly annealed at 200∘C for a couple of hours prior to ARPES. The synchrotron allows to focus a wide range of photon energies on the sample with a beam spot of 5-10 m in diameter and total energy resolution (beamline and analyzer) of 50 meV. For our experiments, the photon beam impinged with an incident angle of with respect to the surface normal and with a azimuth relative to the direction of the crystal. The polarization was within the plane of incidence (p-polarized). The Fermi level has been determined by ARPES on the Au film of the substrate. The sample temperature is adjusted by changing the liquid He flow rate and measured by a Lake Shore DT-670E-BR Si diode placed at the sample holder. So far, only two temperatures are employed for technical reasons such that ARPES data at temperatures closer to have to be postponed to future experiments.
@@
Computational Details: The calculations were performed in the framework of density functional theory (DFT) using the generalized gradient approximation within the PBE flavor PhysRevLett.77.3865, as implemented in VASP software Kresse1996. The ion–electron interactions were described by the projector augmented wave (PAW) method Holzwarth2001. Plane-wave basis cutoff and centered Monkhorst-Pack Monkhorst1976 k-point grid were set to 550 eV and , respectively. A Gaussian smearing of 0.05 eV was employed for the Brillouin zone (BZ) integration. The interlayer vdW forces were treated within the Grimme scheme using D3 correction doi:10.1063/1.3382344. All of the results were obtained using PBE+U method based on Dudarev’s approach PhysRevB.57.1505, with the effective on-site Hubbard U parameter (, where is fixed to eV) for the 3d orbitals. The position of the atoms and unit cell were fully optimized within the PBE+U approach. To simulate the disordered paramagnetic state, we used a supercell with spin-orbit interaction and randomly set the directions of the magnetic spin moments (in 3D) so that the net moment was zero. Here, we did not optimize the geometric structure, in order to capture only changes in electronic dispersion. Moreover, we used the widely employed band unfolding method to obtain an effective electronic structure Popescu2012. Details on additional calculations given in the Supporting Information are provided in Suppl. Section S13.
Associated Content
@@
All data that support the findings of this study are available in zenodo at DOI 10.5281/zenodo.16162323.
Notes
A preprint of this manuscript has been submitted to arXiv Pestka2025_arxiv.
The authors declare no competing financial interest.
@@
A possible application of insulating magnetic 2D materials, details on sample preparation, experimental and computational methods, adaption of and including dependence of ARPES data, individual orbital projections of bands, additional charge density plots, ARPES and DFT band structure along distinct directions, DFT+U band structure of paramagnetic phase, analysis of ARPES structure at eV, detailed description of the use of ARPES curvature and additional DFT calculations employing HSE and using GW methods.
@@ Acknowledgement
We acknowledge financial support from the German Research Foundation (DFG) via the project Mo 858/19-1 and Germany’s Excellence Strategy Cluster of Excellence: Matter and Light for Quantum Computing(ML4Q) EXC 2004/1 (project Nr. 390534769) from the German Ministry of Education and Research (Project 05K2022 -ioARPES). B.B. acknowledges the support of Humboldt Research Fellowship funded by the Alexander von Humboldt Foundation. M.B. acknowledges financial support from the National Science Centre (NCN), Poland [2024/53/B/ST3/04258]. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2024/017490. K.W. acknowledges support by the National Science Centre in Poland under the project 2024/55/B/ST3/03144. A.K.B. and E.L. were supported by the European Commission via the Marie-Sklodowska Curie action Phonsi (H2020-MSCA-ITN-642656).